f(x,y)=tanx².secx²(x+6x+2xy). Compute partial derivatives of f with respect to x, y.


1
Expert's answer
2020-02-07T11:11:58-0500

The function is given by:

f(x,y) = tan (x2) sec(x2) (x+6x+2xy)

= tan (x2) sec(x2) (7x+2xy)

Now:

The partial derivative of f(x,y) with respect to x is:

fx = tan(x2)sec(x2) (7 + 2y) + (7x + 2xy)[ sec (x2) (2x sec2 (x2)) + tan (x2) (2x sec(x2) tan(x2 ))] [ By using Product Rule of Derivative ]

= tan(x2)sec(x2) (7 + 2y) + x(7 + 2y)[ 2x sec3 (x2) + 2x tan2 (x2) sec(x2) ]

= (7 + 2y) sec(x2) ( tan(x2) + 2x2 sec2 (x2) + 2x2 tan2 (x2) )

And:

The partial derivative of f(x,y) with respect to y is:

fy = tan(x2)sec(x2) (0 + 2x) + (7x + 2xy) (0) [ By using Product Rule of Derivative and the partial derivative of tan(x2)sec(x2) with respect to y is zero. ]

= 2x tan(x2)sec(x2)